Development of a new deluxe version of a particular software product is being considered. The activities necessary for the completion of this are listed in the table below (Time in weeks).

Activity

Immediate

Predecessor

a

m

b

A

--

2

5

8

B

A

3

6

9

C

A

4

7

10

D

B

2

5

14

E

D

3

3

3

F

C

6

8

10

G

E, F

1

1

1

H

C

6

10

14

I

G, H

3

4

5

(a) If you wish to reduce the time required to complete this project by two weeks, which activity(ies) should be crashed and why?
Answer #1

We will select activities = A, C, H, I

EXPLANATION:

In order to find the expected time, we will use the following formula

T = (a + 4m + b) / 6

A. (2 + (4 * 5 ) + 8) / 6 = 5

B. (3 + (4 * 6 ) + 9) / 6 = 6

C. (4 + (4 * 7 ) + 10) / 6 = 7

D. (2 + (4 * 5 ) + 14) / 6 = 6

E. (3 + (4 * 3 ) + 3) / 6 = 3

F. (6 + (4 * 8 ) + 10) / 6 = 8

G. (1 + (4 * 1 ) + 1) / 6 = 1

H. (6 + (4 * 10 ) + 14) / 6 = 10

I. (3 + (4 * 4 ) + 5) / 6 = 4

USING THE EXPECTED TIME, WE CAN FIND THE CRITICAL PATH:



ACTIVITY

DURATION

PRE 1

PRE 2

ES

EF

LS

LF

SLACK

A

5

0

5

0

5

0

B

6

A

5

11

6

12

1

C

7

A

5

12

5

12

0

D

6

B

11

17

12

18

1

E

3

D

17

20

18

21

1

F

8

C

12

20

13

21

1

G

1

E

F

20

21

21

22

1

H

10

C

12

22

12

22

0

I

4

G

H

22

26

22

26

0


The ES, EF values are calculated using a forward pass where the ES of the next activity is the maximum EF of all the predecessor activities.

The LS, LF values are calculated using a backward pass where the LS of the next activity is the minimum LF of all the predecessor activities.

The critical path represents all the activities in a particular chain which have the largest duration and 0 slack for all activities, and due to this fact, any delay in the critical activities can cause a delay for the entire project.


A - C - H - I = 26

If we have to crash an activity to reduce the project time, the critical activities would be selected in order to bring about a decrease in the project time and therefore, for crashing, we will select activities = A, C, H, I